Suppose that the matrix A A has the following eigenvalues and eigenvectors:(1 point) Suppose that the matrix A has the following eigenvalues and eigenvectors: 2 = 2i with v1 = 2 - 5i and - 12 = -2i wi
Answer #1

= Aro > () The eigenvalues and eigenvectors of A are as follows: 4 d=20 with = 2 -5i and dq=-zi with ů = [216:] We solve (1)to the system Hence for the linear the general solution system rar be x(+) ar tor2 1 = sin (2+) + 0+ e cos (27) [] J - (18 (A. In eigenvalue / eigenvector form: Oit * (*) sin (2) 4 -2 - ([ Cos (27) @ 5 y (+) cos(2+) 0.t e ta sin (24) + ([11 I using

Please give a thumbs up if you are benefited.